OpenStudy (anonymous):

Evaluate each polynomial for x = -1 and y = 3. -7x + x2y

OpenStudy (hoblos):

"x2y" is that x*2*y or (x^2)*y ??

OpenStudy (anonymous):

its x2y

OpenStudy (anonymous):

\[-7x + x^2y\]

OpenStudy (anonymous):

sooo easy -7(-1) + (-1)2(3) 7+-6= is what

OpenStudy (anonymous):

@ZhangYan its actually: \[-7(-1) + (-1)^2(3) = 7 + 3\] -1^2 =1

OpenStudy (anonymous):

Which is obviously 10

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Your welcome:)